What is the probability that a random teacher from a high school is older than 45 years old, p(x>4)?
1
Expert's answer
2020-06-03T18:55:57-0400

Probability of "p(x>4 ) = 0.5 - p(x) = 0.9999-0.5 = 0.4999 = 0.5 = \frac {1}{2}"